If you are not prepaired for a test and you have 3 true or false questions what is the probablity of getting 2 out of the 3 correct?

Answer by stanbon(75887)

If you are not prepaired for a test and you have 3 true or false questions what is the probablity of getting 2 out of the 3 correct?
-----
Binomial Problem with n = 3 and p(correct) = 1/2
---
P(x = 2 or 3) = P(2) + P(3)
= 3C2(1/2)^3 + 3C3(1/2)^2
----
= (3+1)(1/2)^3
----
= 4/8
---
= 1/2
============
Cheers,
Stan H.
==============